OpenStudy (baldymcgee6):

Given a line of equation of y=mx+b , find the equations of the lines parallel and perpendicular to it and containing, respectively, the point: a) (1,2) b) (xo,yo) c) (a,0)

OpenStudy (baldymcgee6):

how would we find "b"?

OpenStudy (anonymous):

you assume m and b to be constants. You need to find 2 new lines, one of which is parallel to the given line and one of them is perpendicular to the given line. The only condition, these two lines must pass through the given point.

OpenStudy (anonymous):

Line in the question: \[ y = mx + b \] Line parallel to it: \[ y = mx + c \] Line perpendicular to it: \[ y = \frac{-x}{m} + d \] Find values of c and d in terms of m such that they satisfy the given points.

OpenStudy (anonymous):

(x0 ,y0 )......is the specific point on the two line .So for perpendicular line condition is that the product of slopes of line must be -1. And for parallel lines there slope must be equal but they differ in terms of constant (y intercept) i.e. for given line is b . so atlast perpendicular line must have slope -1/m and parallel line must hav slope same as that of given line i.e. m. A nd finally both perpendicular and parallel line differ from given line also by constants say for perpendicular line it is 'z' and for parallel it is 'k'